OpenStudy (anonymous):

What is the slope of the line through the points (5, 2) and (5, -3)

OpenStudy (zehanz):

The slope can be calculated if you have two points by: \(\dfrac{\Delta y}{\Delta x}\). \(\Delta y\) is the difference of the y-coordinates of the points, \(\Delta x\) is the difference of the x-coordinates of the points. Put the values you have here in the formula and see what happens...(you have been warned)

OpenStudy (anonymous):

which is y and x

OpenStudy (zehanz):

x is always the first number, y the second.

OpenStudy (zehanz):

(x, y)

OpenStudy (anonymous):

i thought it was m= y2 - y1/ X2 - X1

OpenStudy (anonymous):

M = -3 - 2/ 5 - 5

OpenStudy (zehanz):

The formula you named is the same as what I said: (-3-2)/(5-5). Brackets are needed to avoid confusion... What about the result of this calculation?

OpenStudy (anonymous):

-5/0

OpenStudy (anonymous):

-5 correct

OpenStudy (zehanz):

Not correct, I'm afraid. Division by 0 is impossible!

OpenStudy (anonymous):

so it is undefined

OpenStudy (zehanz):

It is, as it is with all lines that are vertical...

OpenStudy (anonymous):

so i would put x = -5 ?